A 5.0m long ladder who weighs 180N rests against a vertical wall, making an angle of 53 degrees with the ground. When a person who weighs 800N climbs this uniform ladder and stops one-third of the way up the ladder. Find the normal and frictional forces on the ladder at its base.


1
Expert's answer
2021-11-29T11:44:06-0500

"\sum F_y=0:\ N_A=980\to N_A=980\ (N)" . Answer


"\sum F_x=0:\ -N_B+F_{fr\ A}=0\to F_{fr\ A}=N_B"


"\sum M_A=0:\ 180\cdot1.5+800\cdot2-4N_B=0\to N_B=268\ (N)"


"F_{fr\ A}=N_B=268\ (N)" . Answer